# Stabilization of codimension of persistence barcodes

### v3

@@ 2. Important classes of examples: persistent homology
A filtration $Z$ of topological spaces indexed by $S$ is a collection of topological spaces $(Z_{s}~{}|~{}s\in S)$ along with inclusions $Z_{s}\subseteq Z_{t}$ whenever $s\leq t$. For any non-negative integer $d$, we can set $V_{s}=H_{d}(Z_{s},\mathbb{K})$ to be the degree $d$ simplicial homology of the space $Z_{s}$ with coefficients in $\mathbb{K}$. Further, whenever $s\leq t$, we obtain induced mappings in degree $d$ homology $f_{st}:V_{s}\to V_{t}$ from the inclusions $Z_{s}\subseteq Z_{t}$; see e.g., [18, Section 2.1] as a general reference on simplicial homology. The $S$-persistence module so obtained from this $(V_{s})$ and this $(f_{st})$ is called the (degree $d$) persistent homology (PH) of the filtration and we will denote it by $V=\operatorname{PH}_{d}(Z)$; see e.g., [17] as a general reference on persistent homology.
@@
For each $s\in S$, the Čech complex is the simplicial complex having a $k$-simplex with vertices $x_{i_{0}},\ldots,x_{i_{k}}$ whenever $\bigcap_{j=0}^{k}B_{s}(x_{i_{j}})\neq\emptyset$. The nerve theorem [5] implies that the Čech complex is homotopy equivalent to the union $C_{s}=\bigcup_{x\in X}B_{s}(x)$; see e.g., [17, Section 1.3]. The filtration given by the inclusions $C_{s}\subseteq C_{t}$ is called the Čech filtration. Let us denote this filtration by $\mathrm{\check{C}}(X,M)$, and again, we obtain an $S$-persistence module $\operatorname{PH}_{d}(\mathrm{\check{C}}(X,M))$.
@@
The VR filtration is a computationally efficient alternative to the Čech filtration since only the pairwise distances between the points need be computed. Fortunately, the VR filtration also provides a controlled approximation to the Čech filtration; see [12, Theorem 2.5] for a precise formulation.
@@ 3.1. Quiver generalities
where $Q_{0}=\{1,2,\ldots,n\}$ and each arrow in $Q_{1}$ has the form $(i)\to(i+1)$. The name is justified since all arrows point in the same direction; i.e., are equioriented, and readers familiar with representation theory of Lie algebras will recognize that these quivers are orientations of the type-A Dynkin diagram [19, Section 11].
@@ 3.3. Isoclasses and orbits
It is a classical result of quiver representation theory that two representations are isomorphic (in the category of quiver representations) if and only if they live in the same $G_{\alpha}$-orbit in $\operatorname{Rep}$ [13, Corollary 1.3.1]. Accordingly, we take a $G_{\alpha}$-orbit in $\operatorname{Rep}$ as the definition of an isomorphism class, aka isoclass, for quiver representations.
@@ 3.4. Codimension of isoclasses
Organize the persistent Betti numbers in a triangular rank array [7, Section 2].
@@ 4. A new formula for the codimension of an isoclass
For any index set $S$, given an interval $I\subseteq S$ with birth $b$ and death $d$, we will write $I=\langle b,d\rangle$. We adopt the convention, here and throughout the paper, that angle brackets in our interval notation can be interpreted either as inclusion or exclusion of the endpoints. The results that follow are independent of this choice. Further, recall that we will use the words “bar” and “interval” interchangeably. Suppose that $I=\langle b,d\rangle$ and $J=\langle\beta,\delta\rangle$. Without loss of generality, assume that $b\leq\beta$.
@@
In the quiver representation literature, the barcode is called a lacing diagram. In a seminal 2005 paper [21], Knutson–Miller–Shimozono introduced the notion of an extended lacing diagram and showed that the codimension of an isoclass could be computed by counting crossing strands in a minimal extension; we will not define the minimality condition in the present paper, but see [21, Theorem 3.8] and [9, Section 2] for details. While the Knutson–Miller–Shimozono formulation is combinatorial, our Theorem 4.2 requires only the knowledge of the birth and death times to compute the codimension, and not any knowledge of a special extension of the barcode. Nevertheless, it would be interesting to compare our Theorem 4.2 with the construction of a minimal extended lacing diagram; however, this setting is too general for the TDA framework and investigating it is beyond the scope of the current paper.
@@ 5. Persistence modules and Qcodim
Let $V$ denote a pointwise finite-dimensional $S$-persistence module. The work of Crawley-Boevey [11, Theorem 1.1], which we have recalled in the present paper as Theorem 1.1 of the introduction, implies that
@@ 7.1. Interactions and restrictions of pointbars
Recall that a pointbar in $S$ (respectively $S^{\prime}$) is a bar $I$ such that $b(I)=a=d(I)$ for some $a\in S$ (resp. $a\in S^{\prime})$.
@@
Indeed, denote the pointbar by $\langle a,a\rangle$ and the other bar (possibly a pointbar, or not) by $\langle b,d\rangle$. If the pair is interlaced, we must either have $a<b\leq a$ or $a\leq d<a$ which are both impossible. ∎
@@
Suppose that $I=\langle a,a\rangle$ is a pointbar for $V$ which interacts with a bar $J$ in the barcode for $V$. Then we must have that $a\in S^{\prime}$, $I^{\prime}=\langle a,a\rangle$ is a pointbar for $V^{\prime}$, and $I^{\prime}$ interacts with a corresponding bar $J^{\prime}$ in the barcode for $V^{\prime}$. Moreover, the bars $I$ and $I^{\prime}$ have the same multiplicity in their respective barcodes, and the bars $J$ and $J^{\prime}$ have the same multiplicity in their respective barcodes.
@@
Lemma 7.1 ensures that $I$ and $J$ must interact as a non-separated pair. Write $J=\langle b,d\rangle$ where it is possible that $b=d$ and so $J$ is also a pointbar, and first consider the case that $a<b$. Since $I$ and $J$ are non-separated, it must be the case that there are no elements of $S$ between $a$ and $b$. Since $h\geq 2$, we must have that $\langle a,b\rangle\cap S^{\prime}=\{a,b\}$. Taken together, the previous two statements imply that $a,b\in S^{\prime}$ and there are no elements of $S^{\prime}$ between $a$ and $b$. Hence, $I^{\prime}=\langle a,a\rangle$ is a bar for $V^{\prime}$ and there must exist a bar $J^{\prime}=\langle b,d^{\prime}\rangle$ with $b\leq d^{\prime}\leq d$ with which it interacts. The case $d<a$ is similar. The statement on multiplicities follows from the fact that $V^{\prime}=V|S^{\prime}$. ∎
@@
Suppose that $I=\langle a,a\rangle$ is a pointbar for $V$ which does not interact with any other bars in the barcode for $V$. One of the following must hold:
@@
  • $a\in S^{\prime}$, and thus $I=\langle a,a\rangle$ is a pointbar for $V^{\prime}$ but still does not interact with any other bars for $V^{\prime}$.
@@ 8.1. Qcodim from VR or Čech filtrations
Figure 1. The underlying point cloud has the form $X_{c}$ with $c=9$. Left: 2-skeleton of the VR complex for $s=3$. Right: 2-skeleton of the VR complex for $s=7$.
Figure 2. The $d=0$ (left) and $d=1$ (right) barcodes for $\operatorname{PH}_{d}(\mathrm{VR}(X_{9},\mathbb{R}^{2}))$. The axis labeled “time” denotes values from the indexing set $S=\mathbb{R}_{\geq 0}$. Notice that the $d=0$ barcode is an example of the result of Theorem 8.1. The images were produced using the TDA package library and GUDHI algorithm in the statistical software, R.
@@ 9.2. Distribution of quiver codimension
That is, it would be interesting to consider an extension of the quiver polynomial to general persistence modules which can be approximated by quiver representations in the sense of Definition 6.2. Furthermore, there is a plethora of characteristic classes associated to an isoclass $\Omega$; e.g., the Chern–Schwartz–Macpherson (CSM) and Segre–Schwartz–Macpherson (SSM) classes in cohomology as well as their K-theoretic, and elliptic-cohomology versions; see e.g., [15] for a survey on the current state of the art. These “higher”, non-homogeneous invariants typically recover the quiver polynomial as a specialization to their lowest degree term. The SSM classes may be of particular interest in data analysis, as they can satisfy a normalization property akin to a probability distribution; see, e.g., [14, Remark 8.8].

### v5

@@ 2. Important classes of examples: persistent homology
A filtration $Z$ of topological spaces indexed by $S$ is a collection of topological spaces $(Z_{s}~{}|~{}s\in S)$ along with inclusions $Z_{s}\subseteq Z_{t}$ whenever $s\leq t$. For any non-negative integer $d$, we can set $V_{s}=H_{d}(Z_{s},\mathbb{K})$ to be the degree $d$ simplicial homology of the space $Z_{s}$ with coefficients in $\mathbb{K}$. Further, whenever $s\leq t$, we obtain induced mappings in degree $d$ homology $f_{st}:V_{s}\to V_{t}$ from the inclusions $Z_{s}\subseteq Z_{t}$; see e.g., [19, Section 2.1] as a general reference on simplicial homology. The $S$-persistence module so obtained from this $(V_{s})$ and this $(f_{st})$ is called the (degree $d$) persistent homology (PH) of the filtration and we will denote it by $V=\operatorname{PH}_{d}(Z)$; see e.g., [18] as a general reference on persistent homology.
@@
For each $s\in S$, the Čech complex is the simplicial complex having a $k$-simplex with vertices $x_{i_{0}},\ldots,x_{i_{k}}$ whenever $\bigcap_{j=0}^{k}B_{s}(x_{i_{j}})\neq\emptyset$. The nerve theorem [6] implies that the Čech complex is homotopy equivalent to the union $C_{s}=\bigcup_{x\in X}B_{s}(x)$; see e.g., [18, Section 1.3]. The filtration given by the inclusions $C_{s}\subseteq C_{t}$ is called the Čech filtration. Let us denote this filtration by $\mathrm{\check{C}}(X,M)$, and again, we obtain an $S$-persistence module $\operatorname{PH}_{d}(\mathrm{\check{C}}(X,M))$.
@@
The VR filtration is a computationally efficient alternative to the Čech filtration since only the pairwise distances between the points need be computed. Fortunately, the VR filtration also provides a controlled approximation to the Čech filtration; see [13, Theorem 2.5] for a precise formulation.
@@ 3.1. Quiver generalities
where $Q_{0}=\{1,2,\ldots,n\}$ and each arrow in $Q_{1}$ has the form $(i)\to(i+1)$. The name is justified since all arrows point in the same direction; i.e., are equioriented, and readers familiar with representation theory of Lie algebras will recognize that these quivers are orientations of the type-A Dynkin diagram [20, Section 11].
@@ 3.3. Isoclasses and orbits
It is a classical result of quiver representation theory that two representations are isomorphic (in the category of quiver representations) if and only if they live in the same $G_{\alpha}$-orbit in $\operatorname{Rep}$ [14, Corollary 1.3.1]. Accordingly, we take a $G_{\alpha}$-orbit in $\operatorname{Rep}$ as the definition of an isomorphism class, aka isoclass, for quiver representations.
@@ 3.4. Codimension of isoclasses
Organize the persistent Betti numbers in a triangular rank array [8, Section 2].
@@ 4. A new formula for the codimension of an isoclass
For any index set $S$, given an interval $I\subseteq S$ with birth $b$ and death $d$, we will write $I=\langle b,d\rangle$. We adopt the convention, here and throughout the paper, that angle brackets in our interval notation can be interpreted either as inclusion or exclusion of the endpoints. The results that follow are independent of this choice. However, in Section 7, we will require closed brackets for pointbars $[a,a]$. Further, recall that we will use the words “bar” and “interval” interchangeably. Suppose that $I=\langle b,d\rangle$ and $J=\langle\beta,\delta\rangle$. Without loss of generality, assume that $b\leq\beta$.
@@
In the quiver representation literature, the barcode is called a lacing diagram. In a seminal 2005 paper [22], Knutson–Miller–Shimozono introduced the notion of an extended lacing diagram and showed that the codimension of an isoclass could be computed by counting crossing strands in a minimal extension; we will not define the minimality condition in the present paper, but see [22, Theorem 3.8] and [10, Section 2] for details. While the Knutson–Miller–Shimozono formulation is combinatorial, our Theorem 4.2 requires only the knowledge of the birth and death times to compute the codimension, and not any knowledge of a special extension of the barcode. Nevertheless, it would be interesting to compare our Theorem 4.2 with the construction of a minimal extended lacing diagram; however, this setting is too general for the TDA framework and investigating it is beyond the scope of the current paper.
@@ 5. Persistence modules and Qcodim
Let $V$ denote a pointwise finite-dimensional $S$-persistence module. The work of Crawley-Boevey [12, Theorem 1.1], which we have recalled in the present paper as Theorem 1.1 of the introduction, implies that
@@ 7.1. Interactions and restrictions of pointbars
Recall that a pointbar in $S$ (respectively $S^{\prime}$) is a bar $I$ such that $b(I)=a=d(I)$ for some $a\in S$ (resp. $a\in S^{\prime})$. We will write such a pointbar as $[a]$.
@@
Denote the pointbar by $[a]$ and the other bar (possibly a pointbar, or not) by $\langle b,d\rangle$. If the pair is interlaced, we must either have $a<b\leq a$ or $a\leq d<a$ which are both impossible. ∎
@@
Suppose that $I=[a]$ is a pointbar for $V$ which interacts with a bar $J$ in the barcode for $V$. Then we must have that $a\in S^{\prime}$, $I^{\prime}=[a]$ is a pointbar for $V^{\prime}$, and $I^{\prime}$ interacts with a corresponding bar $J^{\prime}$ in the barcode for $V^{\prime}$. Moreover, the bars $I$ and $I^{\prime}$ have the same multiplicity in their respective barcodes, and the bars $J$ and $J^{\prime}$ have the same multiplicity in their respective barcodes.
@@
Lemma 7.1 ensures that $I$ and $J$ must interact as a non-separated pair. Write $J=\langle b,d\rangle$ where it is possible that $b=d$ and so $J$ is also a pointbar, and first consider the case that $a<b$. Since $I$ and $J$ are non-separated, it must be the case that there are no elements of $S$ between $a$ and $b$. Since $h\geq 2$, we must have that $\langle a,b\rangle\cap S^{\prime}=\{a,b\}$. Taken together, the previous two statements imply that $a,b\in S^{\prime}$ and there are no elements of $S^{\prime}$ between $a$ and $b$. Hence, $I^{\prime}=[a]$ is a bar for $V^{\prime}$ and there must exist a bar $J^{\prime}=\langle b,d^{\prime}\rangle$ with $b\leq d^{\prime}\leq d$ with which it interacts. The case $d<a$ is similar. The statement on multiplicities follows from the fact that $V^{\prime}=V|S^{\prime}$. ∎
@@
Suppose that $I=[a]$ is a pointbar for $V$ which does not interact with any other bars in the barcode for $V$. One of the following must hold:
@@
  • $a\in S^{\prime}$, and thus $I=[a]$ is a pointbar for $V^{\prime}$ but still does not interact with any other bars for $V^{\prime}$.
@@ 8.1. Qcodim from VR or Čech filtrations
Figure 1. The underlying point cloud has the form $X_{c}$ with $c=8$. Top: 2-skeleton of the VR complex for $s=4.5$. Bottom: 2-skeleton of the VR complex for $s=8$. Images created with the software [4].
Figure 2. The $d=0$ (left) and $d=1$ (right) barcodes for $\operatorname{PH}_{d}(\mathrm{VR}(X_{8},\mathbb{R}^{2}))$ with $X_{8}$ as in Figure 1. The axis labeled “time” denotes values from the indexing set $S=\mathbb{R}_{\geq 0}$. Notice that the $d=0$ barcode is an example of the result of Theorem 8.1. The images were produced using the TDA package library and GUDHI algorithm in the statistical software, R.
@@ 9.2. Distribution of quiver codimension
That is, it would be interesting to consider an extension of the quiver polynomial to general persistence modules which can be approximated by quiver representations in the sense of Definition 6.2. Furthermore, there is a plethora of characteristic classes associated to an isoclass $\Omega$; e.g., the Chern–Schwartz–Macpherson (CSM) and Segre–Schwartz–Macpherson (SSM) classes in cohomology as well as their K-theoretic, and elliptic-cohomology versions; see e.g., [16] for a survey on the current state of the art. These “higher”, non-homogeneous invariants typically recover the quiver polynomial as a specialization to their lowest degree term. The SSM classes may be of particular interest in data analysis, as they can satisfy a normalization property akin to a probability distribution; see, e.g., [15, Remark 8.8].